Find the variables and the measure of each angle.
two lines cross each other like an X.
the top is (x+y+5) degrees
the one on the right is (2x) degrees
and the one to the left is (y) degrees
I have tried over and over to solve this problem but I don't know how to solve it with two variables.

Answer by Simnepi(216)

When two straight lines cross opposite angles are equal.
From your description the angle y and the angle 2x are opposite each other. Therefore they are the same.
So now we know that y = 2x.
This means that the top angle in your diagram can be written
(x+2x+5) degrees
This is the same as 3x+5.
Look at your diagram and cover the left and bottom angles with a piece of paper.
You should see that the top angle and the angle on the right are on a straight line. you should know that angles on a line add up to 180 degrees.
So now we can say that
3x+5 + 2x = 180
which means that
5x + 5 = 180
therefore
5x = 175
and so
x = 35 degrees
Now you can see that
y = 70 degrees (2 times 35)(the same as the 2x angle, remember they're equal when they're opposite)
and the top angle is
35 + 70 + 5 = 110 degrees
The bottom angle is the same as the top (they're opposite angles too!)